OpenStudy (anonymous):

Maths Questions:Simultaneous Equations 1. 4x-3y=14 2x-2y=-7 2. 3x+2y=11 2x-5y=1 3. 5x + 4y = 5 3-x5y=-34 4. 4x-3y=5 2x+2y= -1

OpenStudy (ghazi):

which one first?

OpenStudy (anonymous):

any :)

OpenStudy (ghazi):

cool , well lets begin with first one 4x-3y=14 2x-2y=-7 your equation becomes 4x-3y=14 4x-4y=-14 , i have multiplied equation 2 by 2, now subtract equation one from 2 y=28 and put y=28 in any equation to get x hope you can do it now :)

OpenStudy (anonymous):

thanks :)

OpenStudy (ghazi):

similarly you can do the rest :)

OpenStudy (ghazi):

change your second equation of question 3 it should be 3x-5y not 3-x5y ;)

OpenStudy (anonymous):

it says the answers to the first one are 0.5 and -4? how does that work then o_o

OpenStudy (ghazi):

wait a sec let me see again

OpenStudy (ghazi):

answer is correct because it satisfies the equation but the one i have solved shows no mistake either , how is this possible

OpenStudy (ghazi):

i got x=24.5 and y=28 which is also correct

OpenStudy (ghazi):

let me solve the another one , and see what can i get :(

OpenStudy (anonymous):

i agree 0_0

OpenStudy (ghazi):

4x-3y=5 (2x+2y=-1)*2 4x-3y=5 4x+4y=-2 , again subtract equation 1 from 2 7y=-7 y=-1 and from this we can get x 4x-3*-1=5 4x=5+3 4x=8 x=2 so finally i got x=2 and y=-1 , now what does your answer say?

OpenStudy (anonymous):

It said exactly the same as what you got, i think there are multiple ways of answering the question , Thanks :)

OpenStudy (ghazi):

i guess so ;) you're welcome :)